### v3

# GauSSmart: Enhanced 3D Reconstruction through 2D Foundation Models and Geometric Filtering

Abstract
Scene reconstruction has emerged as a central challenge in computer vision, with approaches such as Neural Radiance Fields (NeRF) and Gaussian Splatting achieving remarkable progress. While Gaussian Splatting demonstrates strong performance on large-scale datasets, it often struggles to capture fine details or maintain realism in regions with sparse coverage, largely due to the inherent limitations of sparse 3D training data.
In this work, we propose GauSSmart, a hybrid method that effectively bridges 2D foundational models and 3D Gaussian Splatting reconstruction. Our approach integrates established 2D computer vision techniques, including convex filtering and semantic feature supervision from foundational models such as DINO, to enhance Gaussian-based scene reconstruction. By leveraging 2D segmentation priors and high-dimensional feature embeddings, our method guides the densification and refinement of Gaussian splats, improving coverage in underrepresented areas and preserving intricate structural details.
We validate our approach across three datasets, where GauSSmart consistently outperforms existing Gaussian Splatting in the majority of evaluated scenes. Our results demonstrate the significant potential of hybrid 2D-3D approaches, highlighting how the thoughtful combination of 2D foundational models with 3D reconstruction pipelines can overcome the limitations inherent in either approach alone. Code is available at: https://github.com/alevalve/gaussmart
1 Introduction
Figure 1: Our 3D reconstruction pipeline: An input image undergoes segmentation, which enhances the 2D Gaussian splats, leading to a more geometric 3D mesh.
Three-dimensional (3D) reconstruction has emerged as a prominent area of research since the introduction of Neural Radiance Fields (NeRF) [mildenhall2020nerf]. More recently, Gaussian Splatting [kerbl2023] has advanced scene reconstruction by exploiting the mathematical properties of 3D Gaussians, which allow for both accurate modeling and efficient rendering. Nevertheless, these methods frequently rely on unstructured point clouds, where uneven density and photogrammetry artifacts often introduce noise, limiting the fidelity of the reconstructed geometry.
In addition, optimization strategies in existing approaches are typically restricted to simple objectives such as L1 reconstruction losses or supervision based on surface normals and depth maps. While these objectives provide a useful training signal, they are insufficient to capture fine-grained, high-frequency details that are critical for realistic scene representation. As a result, reconstructions may lack sharpness and fail to preserve subtle geometric and textural structures present in the original scenes.
Given the spurious conditions of the input point clouds, we first apply a convex outlier removal procedure to eliminate unnecessary points introduced by interpolation errors in COLMAP reconstructions. In addition, we present a framework that leverages 2D foundation models such as DINOv3 [simeoni2025dinov3] and SAM [kirillov2023segment] to guide point cloud densification. Specifically, we increase the density of underrepresented segments by sampling new points according to the covariance structure of the original points within each segment mask. Finally, we introduce a novel loss function built on DINOv3 embeddings, which is differentiable and seamlessly integrates into the Gaussian Splatting pipeline [huang2024gaussian], enabling improved semantic alignment during optimization.
Our contributions are summarized as follows:
  • Convex-guided outlier removal. We propose a convex-based filtering method that removes spurious points arising from photogrammetry and interpolation errors in COLMAP-generated point clouds.
  • Segment-aware point cloud densification. By leveraging SAM-derived segmentation masks, we regularize and densify sparse regions of the point cloud using a covariance-driven sampling strategy that accounts for both local geometry and segment area.
  • Embedding-aligned training. We introduce a lightweight embedding loss based on DINOv3 features, which enforces consistency between rendered Gaussian appearances and segment-level semantics, thereby improving spatial fidelity and object coherence.
2 Related Works
Outlier Removal. Outlier removal is a key pre-processing step for noisy point clouds, with two widely used methods being Statistical Outlier Removal (SOR) and Radius Outlier Removal (ROR), both popularized in the Point Cloud Library [Rusu2011PCL, PCL_SOR_ROR]. SOR computes the mean distance of each point to its $k$-nearest neighbors and rejects those deviating from the global mean–standard deviation range, while ROR removes points with fewer than a threshold number of neighbors within a fixed radius. These lightweight filters are usually combined with robust model-fitting methods such as RANSAC [Fischler1981RANSAC] or Trimmed ICP [Chetverikov2002TICP], surface-fitting and projection approaches like Point Set Surfaces [Alexa2003PSS], graph-based and variational denoisers [Zeng2019GraphDenoise, Bouaziz2013L0], and learning-based frameworks such as PointCleanNet [Rakotosaona2019PointCleanNet].
Novel View Synthesis Neural Radiance Fields (NeRF)[mildenhall2020nerf] model view-dependent color and density using a multilayer perceptron (MLP) to synthesize novel views. Subsequent variants have improved sampling strategies, scalability, and geometric fidelity[barron2021mipnerf, Liu20neurips_sparse_nerf, Lindell20arxiv_AutoInt, barron2022mipnerf360, yariv2023bakedsdf, cao2024lightning, oechsle2021unisurf, wang2021neus, yariv2021volume, meng_2023_neat, garbin2021fastnerfhighfidelityneuralrendering, lindell2021autoint]. Despite these advances, conventional NeRF models remain computationally expensive and often fail to accurately reconstruct large-scale or texture-poor regions. More recently, scene representations based on projected Gaussian primitives have enabled fast, high-quality rendering while remaining scalable to large scenes [kerbl2023, huang2024gaussian, choi2024meshgsadaptivemeshalignedgaussian, Dalal_2024, guédon2023sugarsurfacealignedgaussiansplatting]. In particular, 2D Gaussian Splatting [huang2024gaussian] improves geometric alignment by orienting primitives along local surface tangents, typically guided by estimated normals. However, coverage gaps may still arise when the initialization is suboptimal or when point density fails to adequately represent fine-scale structures.
2D Foundational Models. Promptable segmentation models such as SAM [kirillov2023segment] provide strong per-object cues that have been successfully applied in both vision and medical reconstruction [morshuis2024segmentationguidedmrireconstructionmeaningfully, Ma_2024, li2024anatomaskenhancingmedicalimage, zhou2025distillationlearningguidedimage, shang2024jointsegmentationimagereconstruction]. For 3D Gaussians, SAGA [cen2025segment3dgaussians] extends SAM-like capabilities but remains limited when handling fine-scale objects. In parallel, semantic backbones such as DINOv2 [oquab2023dinov2] and DINOv3 [simeoni2025dinov3] provide dense, transferable feature embeddings that capture both local detail and global context. In particular, DINOv3 offers semantically consistent representations with strong invariance to viewpoint and appearance changes, making it especially well-suited for embedding-level supervision. This motivates our use of a DINOv3-based loss to complement pixel-domain objectives and enforce semantic alignment in the Gaussian Splatting pipeline.
Point-based representations. Point clouds are flexible, permutation-invariant 3D carriers for reconstruction [yu2024pointdreamer, deng2020cvxnet, melaskyriazi2023pc2projectionconditionedpointcloud] and for tasks like 2D projection/detection [6923303] and point-level segmentation [qi2017pointnetdeeplearningpoint]. Point-based rendering learns to rasterize point features to images [rakhimov2022npbgacceleratingneuralpointbased, zhu2024rpbgrobustneuralpointbased], while self-supervised objectives densify or regularize point sets [Xiao_2023]. Broader advances span reconstruction [yu2022pointbertpretraining3dpoint, pang2022maskedautoencoderspointcloud, zhang2022pointm2aemultiscalemaskedautoencoders, dai2021can3dfast3dmedical, xu2023stop3dtargetreconstruction, sun20233dfusionrealtime3dobject, pan2023multidimensionunifiedswintransformer], completion [yang2021sanetshuffleattentiondeep, xie2018learningdescriptornetworks3d, zhang2021viewguidedpointcloudcompletion, yu2021pointrdiversepointcloud], and point operators [thomas2019kpconvflexibledeformableconvolution, liu2019relationshapeconvolutionalneuralnetwork, Guo_2021].
3 Method
Figure 2: Overview of the GauSSmart pipeline, illustrating the stages of camera clustering, Convex Hull–based outlier removal, point cloud upsampling, and the integration of a DINOv3-based loss.
3.1 Convex Outlier Removal
Point clouds generated by COLMAP often contain hundreds of spurious points that do not contribute meaningfully to the scene of interest. These outliers typically arise from triangulation errors, introducing noise into the reconstruction. In figure 3 we can see external points that do not represent the object of interest. To address this, traditional methods such as Radius Outlier Removal and Statistical Outlier Removal are widely used. These approaches are considered the classical techniques for denoising point cloud representations.
Figure 3: Visualization of a DTU scene point cloud from different viewpoints, showing point cloud outliers even on closed environments
However, most of these methods depend heavily on hyperparameters that are difficult to tune consistently across different scenes, which makes the cleaning process cumbersome. Moreover, they often suffer from limitations: either removing more points than necessary, or failing to eliminate certain outliers due to sparsity issues. Since these techniques generally rely on kkNN-based strategies that require a minimum number of neighboring points, so they struggle in scenarios with irregular density.
Figure 4: Visualization of the convex hull removal process. Left: the original point cloud containing outliers projected outside the valid viewing cone. Right: the refined point cloud after removing convex hull outliers, which eliminates spurious points and preserves only the consistent geometry.
To overcome these challenges, we employed a convex hull–based outlier removal method, introduced in [valverde2025convex], which leverages geometric properties to more reliably filter noise while preserving the integrity of the underlying structure. Unlike previous approaches, this method does not rely on local neighborhoods; instead, it evaluates each point independently and removes it if its distance to the convex hull exceeds a threshold.
3.2 Camera Clustering and Image Selection
Large multi-view datasets often contain hundreds of images, making segmentation across all views computationally prohibitive. Drawing inspiration from camera-parameter–driven view-reduction and calibration strategies [mauro2013overlapping, khan2019fast, yan2024livelearncontinualaction, article, waleed2024cameracalibrationgeometricconstraints], we propose a clustering-based approach to select a compact, geometry-aware subset of views.
Our method extracts camera centers and forward axes from explicit camera-to-world matrices $\mathrm{c2w}$. To ensure robustness to scene scale and anisotropy, we normalize camera positions by subtracting the mean and dividing by the per-axis standard deviation. We then apply $k$-means clustering to these normalized positions and determine the optimal number of clusters $k$ by evaluating candidate values in the range $k\in[\text{min},\min(15,\lfloor N/2\rfloor)]$.
The selection criterion maximizes a composite objective function that balances two key factors: coverage, measured as the average of spatial spread and inter-view angular diversity within each cluster, and compactness, quantified as negative inertia normalized by the data norm. The objective function is formulated as:
$\text{score}(k)\;=\;\alpha\,\underbrace{\text{coverage}(k)}_{\text{spread}+\text{angle div.}}\;+\;\beta\,\underbrace{\text{compactness}(k)}_{-\text{inertia}/\|X\|}.$ (1)
After clustering, we select one representative camera per cluster by combining proximity to the de-normalized cluster center with angular uniqueness within the cluster. Each candidate view is scored as the average of two normalized terms: (i) a distance-based component $1/(1+\text{dist})$, and (ii) the mean angular separation to other cluster members, linearly scaled to the interval $[0,1]$. The highest-scoring view from each cluster is selected as the representative.
This approach preserves both spatial coverage and viewpoint diversity while substantially reducing the computational burden by limiting the number of images processed during segmentation. We apply SAM2 large model  [ravi2024sam2segmentimages] to the selected representative views, generating segmentation masks that include binary regions, bounding boxes, and area measurements. These masks are subsequently used to assign stable segment identifiers to 3D points for downstream processing tasks.
3.3 Segmentation and Multi-View Projection
After applying the segmentation process to the respective images to obtain per-view masks, each 3D point is assigned to image segments using a standard projection pipeline  [6923303]: we normalize coordinates to the dataset scale, transform points into the camera frame using the extrinsics (rotation and translation), and project onto the image plane with the intrinsics, followed by perspective division to obtain pixel locations. For some specific scenarios is required to apply small dataset-specific adjustments (coordinate conventions, calibration differences) and verify projections with visibility checks to ensure accurate point-to-pixel associations across views.
$\lambda\begin{bmatrix}u\\ v\\ 1\end{bmatrix}=\mathbf{K}\,[\mathbf{R}\;\mathbf{t}]\begin{bmatrix}\mathbf{P}\\ 1\end{bmatrix},\qquad(u,v)=\left(\frac{x}{z},\,\frac{y}{z}\right)$
To achieve consistent segment identities across views, we construct a global correspondence map. Segments corresponding to the same object are linked when their 3D point assignments exhibit high normalized overlap. This overlap-based matching yields stable global IDs and robust segment coherence on the fused point cloud, mitigating variability in per-view masks and supporting reliable downstream reconstruction. Importantly, since each 3D point has a unique spatial location, we retain its assignment to the first segment in which it appears. While the same point may project into multiple views, these additional projections provide redundant evidence but no new positional information. Consequently, we discard duplicate assignments, as a point cannot physically belong to multiple objects. This ensures that each point maintains a single, unambiguous label, preserving geometric consistency and preventing fusion artifacts.
3.4 Mask-Area-Guided Point Cloud Enhancement
Previous research on point cloud upsampling has demonstrated that increasing point density significantly improves reconstruction quality [yu2018punetpointcloudupsampling, li2019puganpointcloudupsampling, li2021pointcloudupsamplingdisentangled, rong2024repkpu]. However, uniform densification fails to account for varying importance and visibility across different semantic regions within the scene. Our approach leverages segmentation mask areas from representative views to guide targeted point augmentation. Rather than applying uniform thresholds, we utilize mask area information to estimate appropriate point density for each segment, ensuring visually prominent regions receive adequate geometric representation. For each segment $s_{i}$ with mask area $A_{i}$, we determine the target number of points based on the square root relationship between area and required sampling density:
$n_{\text{target}}=\max\left(\lfloor\sqrt{A_{i}}\cdot\gamma\rfloor,n_{\min}\right),$ (2)
where $\gamma=0.1$ is an empirically determined scaling factor and $n_{\min}=10$ ensures minimum representation for small segments. For segments where the current point count $|P_{s_{i}}|$ falls below the target, we compute the augmentation requirement:
$n_{\text{add}}=n_{\text{target}}-|P_{s_{i}}|,$ (3)
subject to $n_{\text{add}}>0$ and the requirement that segments contain at least $n$ existing points to provide sufficient geometric context. The augmentation process generates new points through statistical sampling from existing segment geometry. New point positions $\mathbf{p}_{\text{new}}$ are sampled from the existing distribution with Gaussian noise for geometric diversity:
$\mathbf{p}_{\text{new}}=\mathbf{p}_{\text{base}}+\boldsymbol{\epsilon},\quad\mathbf{p}_{\text{base}}\sim P_{s_{i}},\quad\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I}),$ (4)
where $P_{s_{i}}$ represents existing points in segment $s_{i}$, and $\sigma$ is adaptively determined based on local point density. Color attributes for augmented points are derived through interpolation from existing segment points, preserving visual consistency across semantic regions.
3.5 Feature Embedding Loss
We build upon recent advances in semantic feature–based supervision [gong2025dinoslamdinoinformedrgbdslam] by introducing a loss term grounded in DINOv3 feature embeddings [simeoni2025dinov3]. Traditional photometric losses such as $\ell_{1}$ or SSIM capture only low-level pixel correspondences and thus fail to encode high-level semantic cues. To address this limitation, we incorporate a DINO-based supervision signal that enforces semantic consistency between ground-truth and rendered images. Prior works typically apply magnitude-based objectives such as $\ell_{1}$ distance directly on feature vectors, which conflates intensity differences with semantic correspondence. In contrast, we propose a cosine similarity formulation that emphasizes the angular relationship between embeddings, thereby isolating semantic correspondence while remaining invariant to feature magnitude scaling. Formally, given embeddings $f_{\text{gt}},f_{r}\in\mathbb{R}^{d}$ from ground-truth image $I_{\text{gt}}$ and rendered image $I_{r}$, we define
$\cos(f_{\text{gt}},f_{r})=\frac{f_{\text{gt}}\cdot f_{r}}{\|f_{\text{gt}}\|_{2}\,\|f_{r}\|_{2}},$ (5)
with the DINO loss given by
$L_{\text{DINO}}(f_{\text{gt}},f_{r})=\lambda_{\text{dino}}\cdot\cos(f_{\text{gt}},f_{r}),$ (6)
where $\lambda_{\text{dino}}$ balances the contribution of semantic supervision. For structural fidelity, we retain a photometric objective composed of an $\ell_{1}$ term and a structural dissimilarity penalty (DSSIM):
$L_{\text{photo}}=(1-\lambda_{\text{dssim}})\cdot L_{1}+\lambda_{\text{dssim}}\cdot\big(1-\text{SSIM}(I_{\text{gt}},I_{r})\big).$ (7)
The overall training objective combines the semantic and photometric components:
$L_{\text{total}}=L_{\text{photo}}+L_{\text{DINO}}.$ (8)
DTU
Refer to caption Refer to caption Refer to caption Refer to caption
Mip-NeRF 360
Refer to caption Refer to caption Refer to caption Refer to caption
T&T
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: DINOv3 similarity heatmaps across DTU, NeRF, and TYT scenes, highlighting the regions most attended by the model. Brighter areas correspond to stronger alignment with the global embedding.
This formulation integrates seamlessly into the 2D Gaussian Splatting framework [huang2024gaussian]. The additional DINO term complements standard perceptual losses by aligning rendered views with the structural and semantic characteristics of the reference images. Consequently, each Gaussian primitive is optimized not only for photometric accuracy but also for semantic coherence, yielding reconstructions that better reflect the underlying scene semantics.
4 Experiments
We now present experiments for our GauSSmart approach, evaluating both appearance and geometry reconstruction across the three benchmark datasets used in this study. Specifically, we conduct experiments on DTU [jensen2014large], Mip-NeRF 360 [barron2022mipnerf360], and Tanks and Temples [Knapitsch2017]. Our dataset preparation follows the same strategy introduced in the 2DGS paper [huang2024gaussian].
4.1 Implementation Details
Hardware and Environment. All experiments are conducted on a single NVIDIA RTX A5000 GPU with 24 GB of memory. Our method is implemented in PyTorch 1.12 with CUDA 11.6. Due to hardware differences compared to the original 2DGS implementation (RTX 3090), direct efficiency comparisons are not feasible.
Loss Implementation. We employ DINOv3 with a Vision Transformer base architecture and $16\times 16$ patch size (ViT-B/16) as our feature extractor, selected after systematic evaluation across different DINOv3 variants on the Tanks and Temples dataset. The DINO loss weight is set to $\lambda=0.05$, determined through empirical evaluation over the range $[0.01,0.9]$ and guided by the weighting strategies described in [huang2024gaussian].
Gaussian Point Processing. Our segmentation-aware sampling strategy enforces a minimum threshold of five points per segment during new sample generation. This constraint is critical, as the covariance matrix computation requires sufficient point density to maintain numerical stability and avoid ill-conditioned matrices. Segments falling below this threshold are either merged with neighboring segments or discarded during preprocessing.
Outlier Removal. We adopt the convex-hull–based outlier removal procedure described in [valverde2025convex], applying it during both initialization and refinement to maintain geometric consistency across viewpoints.
Training Configuration. Unless otherwise specified, we train for 30,000 iterations, following the standard 2DGS optimization schedule.
5 Results
Our Gaussmart implementation demonstrates competitive performance across benchmark datasets. Table 1 reports a quantitative comparison against 3D Gaussian Splatting (3DGS)[kerbl2023] and 2D Gaussian Splatting (2DGS)[huang2024gaussian]. In terms of PSNR, our method outperforms 2DGS in four of six scenes, indicating reduced noise and improved fidelity. While 2DGS falls short of 3DGS on the Tanks and Temples dataset, the incorporation of our proposed loss and point cloud upsampling narrows this gap.
[…]
Barn Caterpillar Courthouse Ignatius MeetingRoom Truck mean
PSNR
Gaussmart 27.20 22.83 21.61 21.60 24.90 24.19 23.72
2DGS 27.15 22.78 21.55 21.68 24.90 24.16 23.70
3DGS 27.51 23.38 22.22 21.53 25.19 24.25 24.01
SSIM
Gaussmart 0.843 0.772 0.765 0.766 0.859 0.848 0.809
2DGS 0.843 0.773 0.766 0.767 0.859 0.848 0.810
3DGS 0.852 0.791 0.779 0.776 0.866 0.853 0.820
LPIPS
Gaussmart 0.204 0.245 0.282 0.203 0.209 0.177 0.220
2DGS 0.204 0.243 0.282 0.204 0.208 0.177 0.220
3DGS 0.160 0.190 0.266 0.153 0.141 0.108 0.170
Our method surpasses existing approaches on the DTU dataset in PSNR, demonstrating clear improvements in reconstruction quality. The controlled acquisition conditions—high-quality cameras, uniform lighting, and consistent viewpoints—enable our model to fully exploit the available visual information. Sparse initial point clouds in DTU also benefit from our densification strategy, which introduces additional geometric primitives to capture fine details.
[…]
3DGS 35.76
SuGaR 34.57
2DGS 34.52
Gaussmart (Ours) 36.30
These stable imaging conditions provide an ideal setting for our DINO-based loss, allowing semantic features to remain consistent across views. The synergy between controlled capture, point cloud densification, and semantic feature preservation explains the consistent PSNR gains observed on the DTU evaluation set.
[…]
scan24 scan37 scan40 scan55 scan63 scan65 scan69 scan83 scan97 scan105 scan106 scan110 scan114 scan118 scan122 mean
PSNR
Gaussmart 36.42 32.99 35.57 35.54 38.98 36.15 34.69 35.86 33.68 36.72 37.63 35.96 34.65 39.22 39.07 36.30
SSIM
Gaussmart 0.964 0.962 0.962 0.938 0.975 0.936 0.940 0.921 0.929 0.933 0.941 0.934 0.936 0.941 0.937 0.941
LPIPS
Gaussmart 0.062 0.091 0.093 0.146 0.075 0.157 0.132 0.223 0.177 0.174 0.209 0.225 0.192 0.225 0.239 0.165
[…]
bicycle flowers garden stump treehill room counter kitchen bonsai mean
PSNR
Gaussmart 24.78 21.05 26.96 26.41 22.41 31.51 28.82 31.18 31.94 27.23
2DGS 24.87 21.15 26.95 26.47 22.27 31.06 28.55 30.50 31.52 26.93
3DGS 25.24 21.52 27.41 25.07 22.49 30.63 28.70 30.32 31.98 27.26
SSIM
Gaussmart 0.725 0.570 0.841 0.760 0.622 0.921 0.960 0.925 0.940 0.807
2DGS 0.752 0.588 0.852 0.765 0.627 0.912 0.900 0.919 0.933 0.805
3DGS 0.771 0.605 0.868 0.775 0.638 0.914 0.905 0.922 0.938 0.815
LPIPS
Gaussmart 0.282 0.382 0.150 0.266 0.380 0.206 0.198 0.129 0.188 0.242
2DGS 0.218 0.346 0.115 0.222 0.329 0.223 0.208 0.133 0.214 0.223
3DGS 0.205 0.336 0.103 0.210 0.317 0.220 0.204 0.129 0.205 0.214
In Fig. 6, our results are comparable to those of 3DGS, while exhibiting noticeably less noise than 2DGS. This improvement stems from reducing noise in the underlying point cloud, which often introduces bubble-like artifacts and leads to noisy renderings. Our reconstructions also preserve colors more consistently, confirming that the DINO-based loss captures and retains high-level semantic features. For outdoor scenes, our approach outperforms 2DGS across multiple cases, demonstrating strong generalization beyond indoor environments. This is particularly evident in the treehill scene, where complex structures such as root systems and tree trunks benefit from our feature-aware loss.
For indoor environments, GauSSmart consistently achieves superior performance across all three quantitative metrics, surpassing both 2DGS and 3DGS. This improvement highlights the robustness of our method under controlled lighting conditions, where detailed textures and diverse scene elements are more effectively reconstructed.
GT Ours 2DGS 3DGS
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 6: A side-by-side comparison of results from our method and the baselines.
Qualitative comparisons in Fig. 7 further confirm the benefits of our approach. Structural elements such as the walls and sky regions in the Caterpillar scene are reconstructed more cleanly and stably than with 2DGS or 3DGS. While 2DGS introduces noisy textures and bubble-like artifacts in low-texture areas, Gaussmart suppresses such instabilities, producing smoother surfaces and more natural transitions. Beyond large homogeneous regions, our method also demonstrates sharper preservation of edges and fine object boundaries, as in the courthouse scene.
GT Ours 2DGS 3DGS
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 7: A side-by-side comparison of results from our method and the baselines.
To further assess generalization, we conduct experiments on the DTU dataset, which consists of object-centric scenes captured under controlled conditions. As shown in Table 2, our method outperforms recent baselines, including 2DGS, 3DGS, and SuGaR, achieving a PSNR improvement of nearly one point over the best competing approach. This highlights the ability of our model to produce cleaner and more faithful reconstructions. We attribute these improvements to the synergy between DTU’s multi-view acquisition and our DINO-based loss, which leverages high-level features to enhance geometry-aware consistency.
Per-scene results in Table 3 further support this conclusion. Gaussmart achieves a mean PSNR of 36.30 dB, with several scans (e.g., 63, 118, 122) exceeding 39 dB, demonstrating robustness across diverse geometries.
6 Ablations
Table 5 presents the mean PSNR across ablation configurations on Tanks and Temples.
Table 5: Ablation study results applied on Tanks and Temples Training Scenes (mean PSNR).
Experiment Mean PSNR
DINO/7000 loss 23.634
DINO/3000 loss 23.611
Just PC Improvement 23.611
Hull/PC Improvement 23.637
Full model 23.641
Point cloud improvements alone achieve limited gains, indicating that densification in isolation is insufficient. Combining hull filtering with point cloud improvements increases mean PSNR, showing that geometric outlier removal stabilizes 3D reconstructions. Our complete model, which integrates DINO guidance with hull filtering and point cloud improvements, attains optimal performance, confirming the complementary benefits of semantic loss and geometric refinement.
6.1 Limitations
Despite the promising results, our approach presents certain computational challenges that warrant consideration. The necessity to compute DINO embeddings each time an image passes through the pipeline introduces a significant computational overhead, potentially increasing processing time by up to 1.5x depending on the size and complexity of the scene. Additionally, our approach computes SAM maps for each image during processing, which adds further computational burden.
Through our experiments, we discovered consistent patterns across datasets: DTU scenes consistently utilize 3 images given that the cameras are the same for all objects, while Mip-NeRF 360 exhibits a similar pattern with 15 images. This observation suggests a promising avenue for optimization—storing that could significantly reduce training time, that is the main drawback of the current approach.
7 Conclusions
In this work, we introduced GauSSmart, a novel approach that combines the strengths of 2D foundational models and 3D methods. Our method highlights the potential of integrating established 2D computer vision techniques, including convex filtering and other foundational components, with 3D reconstruction pipelines. Experimental evaluation across three diverse datasets shows that GauSSmart consistently outperforms existing Gaussian Splatting methods in most scenes, supporting our hypothesis that hybrid 2D–3D approaches can effectively leverage the complementary strengths of both domains. Its superior performance stems from the strategic use of robust 2D priors to compensate for the limitations of sparse 3D training data, underscoring a key insight for the field: the scarcity of high-quality 3D data necessitates intelligent integration with rich 2D image representations.
Looking ahead, the recent release of DiNOv3 presents promising oportunities for advancing 2D foundational models in 3D reconstruction tasks. The enhanced capabilities and improved feature representations offered by DiNOv3 could unlock further greater performance gains when integrated into Gaussian Splatting frameworks.
Supplementary Materials
Geometric Results
Table 6: DTU evaluation results using 2DGS pipeline. Metrics are mean distance-to-surface (d2s), mean surface-to-distance (s2d), and overall error.
Scan mean_d2s mean_s2d overall
24 2.6143 1.4773 2.0458
37 1.8802 0.8228 1.3515
40 2.0464 1.6955 1.8709
55 1.6924 0.6944 1.1934
63 2.8613 2.6491 2.7552
65 2.7081 1.8333 2.2707
69 2.0261 1.1439 1.5850
83 2.2375 1.8825 2.0600
97 1.9723 1.8632 1.9178
105 1.9679 1.4783 1.7231
106 2.4727 1.1019 1.7873
110 2.6417 1.2023 1.9220
114 1.5179 0.9364 1.2272
118 1.8501 0.7577 1.3039
122 2.1410 0.8514 1.4962
Table 7: Indoor scenes (DTU, Nerf indoor, Meetingroom).
Category Mean Points Added Mean Segments with New Points
Indoor Scenes 116.08 8.44
Table 8: Outdoor scenes (Tanks except Meetingroom, Nerf outdoor).
Category Mean Points Added Mean Segments with New Points
Outdoor Scenes 88.15 9.36
Camera Clustering
Examples of the images selected by clustering method for a set of scenes from DTU, MipNerf360 and Tanks and Temples
000045.jpg 000050.jpg 000088.jpg 000147.jpg
Refer to caption Refer to caption Refer to caption Refer to caption
000003.jpg 000009.jpg 000037.jpg 000111.jpg
Refer to caption Refer to caption Refer to caption Refer to caption
000112.jpg
Figure 8: Selected clustered camera views from Barn and Caterpillar scenes.
0000.png 0004.png 0045.png
Refer to caption Refer to caption Refer to caption
0000.png 0011.png 0019.png
Refer to caption Refer to caption Refer to caption
Figure 9: Clustered camera views from DTU scans: Scan65 (top row) and Scan122 (bottom row).
Bicycle
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Counter
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 10: Selected clustered camera views — Bicycle (top) and Counter (bottom).